What is the answer for 16-3x=4+x
Found 3 solutions by stanbon, ikleyn, Alan3354:
Answer by stanbon(75887)

What is the answer for 16-3x=4+x
------
4x = 16-4 = 12
x = 3
--------------
Cheers,
Stan H.
-------------

Answer by ikleyn(52788)

.
Read the lesson
    - HOW TO solve a linear equation
in this site.


Answer by Alan3354(69443)

Depends on the question.